OpenStudy (anonymous):

Solve for x : √4x+5 = x

OpenStudy (ganpat):

1st step - square them on both side 2nd step - bring all the terms together 3 rd step - solve the quadratic equation..

OpenStudy (anonymous):

so would it look like this ? √4x+5 = √x

OpenStudy (ganpat):

no no.. 1st tell me, is root (4x) .. or just root (4) ???

OpenStudy (anonymous):

I think just root (4)

OpenStudy (ganpat):

so root of 4 = 2 so no need to square than... now equation becomes 2x +5 = x 2x - x = -5 x = -5... solution...

OpenStudy (anonymous):

Ohhhhhh , I see now .. Thank you ! (:

OpenStudy (ganpat):

:)